# Robustness of topological phases on aperiodic lattices

### v1

@@ 3.1.1. The regular representation
The reduced C*-algebra of an étale groupoid $\mathcal{G}$ is the completion of the groupoid convolution algebra $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$ under the regular representation (cf. [Khoshkam-Skandalis:Regular_representation]), which we recall here. The real groupoid convolution algebra $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$ consists of real-valued, compactly supported continuous functions $f\colon\mathcal{G}\to\mathbb{R}$, equipped with the convolution product and involution
@@
The space $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$ is a pre-Hilbert module over $\mathrm{C}_{0}(\mathcal{G}^{0})$, with right $\mathrm{C}_{0}(\mathcal{G}^{0})$-module structure and inner product (cf. [Bourne-Mesland:Aperiodic]*Section 1.3)
$(f\cdot\varphi)(\gamma)$ :$\mathrel{\vcentcolon=}f(\gamma)\varphi(s(\gamma)),$
$\left\langle f,g\right\rangle(x)$ :$\mathrel{\vcentcolon=}(f^{*}*g)\big{|}_{\mathcal{G}^{0}}(x)=\sum_{\gamma\in r^{-1}(x)}f(\gamma)g(\gamma),$
@@
[…]
which is in bijection with the Delone set $\omega$ via $(\omega-x,-x)\mapsto x$. Thus the localised Hilbert space $\mathcal{H}_{\omega}$ is unitarily isomorphic to $\ell^{2}(\omega)$ via (cf. [Bourne-Mesland:Aperiodic]*Section 4.1):
@@
Lemma 3.5 ([Bourne-Mesland:Aperiodic]*Section 4.1).
@@
When we consider systems with a finite number of internal degrees of freedom inside every lattice site, e.g. spins of electrons, then we must replace the observable C*-algebra by its matrix algebra $\mathbb{M}_{N}\mathrm{C}^{*}(\mathcal{G}_{\Lambda})=\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$. We extend the regular representation of $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ to a representation
$\pi_{\omega}^{N}\mathrel{\vcentcolon=}\pi_{\omega}\otimes e_{N},\quad\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})\to\mathbb{B}(\ell^{2}(\omega,\mathcal{K}))$
where $\mathcal{K}$ is a separable Hilbert space, and $e_{N}\colon\mathbb{M}_{N}(\mathbb{R})\hookrightarrow\mathbb{K}(\mathcal{K})$ is any rank-$N$ corner embedding. We may describe the representation $\pi_{\omega}^{N}$ by an infinite matrix indexed by $x,y\in\omega$, with matrix elements
@@ 3.1.2. Position spectral triple over the groupoid C*-algebra
In order to construct a spectral triple over $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ or $\mathbb{M}_{N}\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, one way is to first construct an unbounded Kasparov $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$-$\mathrm{C}(\Omega_{0})$–module, then take its (unbounded) Kasparov product with a *-homomorphism $\mathrm{C}(\Omega_{0})\to\mathbb{R}$. Such an unbounded Kasparov module was given in [Bourne-Mesland:Aperiodic]*Section 2.3.1 using the “position” operators on the Hilbert $\mathrm{C}(\Omega_{0})$-module $L^{2}(\mathcal{G}_{\Lambda})$:
@@
The map $c_{j}\colon\mathcal{G}_{\Lambda}\to\mathbb{R}$ is an exact groupoid cocycle in the sense of [Mesland:Groupoid_cocycles]*Definition 4.1.2. It follows from the general construction in [Mesland:Groupoid_cocycles]*Theorem 3.2.2 that $\prescript{}{d}{\lambda}_{\Omega_{0}}$ is an unbounded Kasparov module. Following [Bourne-Mesland:Aperiodic], we call $\prescript{}{d}{\lambda}_{\Omega_{0}}$ the bulk cycle.
[…]
@@
(3.10) $\xi_{\omega,N}^{\mathrm{Gpd}}\mathrel{\vcentcolon=}\left(\mathbb{M}_{N}(\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda}))\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\omega,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)$
is a locally compact position spectral triple. It represents a class in
@@
which is the Kasparov product of the following unbounded Kasparov modules and *-homomorphisms:
@@
    exterior product with the imprimitivity bimodule $(\mathbb{M}_{N}(\mathbb{R}),\mathcal{K},0)$, which represnts the identity element of $\operatorname{KK}(\mathbb{R},\mathbb{R})$.
@@
is isomorphic to $\ell^{2}(\omega,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$ via the localised representation at $\omega$. Thus it suffices to check that $c_{j}$ is mapped to ${\scriptstyle\mathsf{X}}_{j}$ under the composition and that (3.10) is a spectral triple. By (3.4), the operator $X_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\operatorname{id}_{\bigwedge\nolimits^{*}\mathbb{R}^{d}}$ acts on $\ell^{2}(\omega,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$ sends $\left\lvert x\right\rangle\otimes v\otimes w$ to $x_{j}\left\lvert x\right\rangle\otimes v\otimes w$, where $x\in\omega$ is considered an element in $\mathbb{R}^{d}$. It is a spectral triple by [Bourne-Mesland:Aperiodic]*Proposition 4.1. Thus (3.10) is a position spectral triple. It is locally compact in the sense of Definition 2.16 because the matrix elements of $\pi^{N}_{\omega}$ belong to $\mathbb{M}_{N}(\mathbb{R})$, hence are compact operators on $\mathcal{K}$. ∎
@@ 3.2. Roe C*-algebra of a Delone set
Let $X$ be a proper metric space. An ample $X$-module or ample module over $X$ is given by a pair $(X,\mathcal{H}_{X})$, where $X$ is a proper metric space, and $\mathcal{H}_{X}$ is a Hilbert space that carries a non-degenerate representation $\varrho\colon\mathrm{C}_{0}(X)\to\mathbb{B}(\mathcal{H}_{X})$, which is ample in the sense that
$\varrho(f)\in\mathbb{K}(\mathcal{H}_{X})$ iff $f=0$.
@@
  • An operator $S\in\mathbb{B}(\mathcal{H}_{X})$ is called locally compact, if for any $f\in\mathrm{C}_{\mathrm{c}}(X)$, $T\varrho(f)$ and $\varrho(f)T$ are compact operators.
  • An operator $S\in\mathbb{B}(\mathcal{H}_{X})$ is called controlled, if $\operatorname{supp}(S)$ is controlled.
@@
Let $X$ be a discrete metric space and $\mathcal{H}_{X}=\ell^{2}(X,\mathcal{K})$ be the standard ample $X$-module. Describe an operator $T\in\mathbb{B}(\mathcal{H}_{X})$ by an infinite matrix $(T_{x,y})_{x,y\in X}$ where
@@
The operator $T$ is locally compact, iff $T_{x,y}\in\mathbb{K}(\mathcal{K})$ for all $x,y\in X$; controlled, iff there exists $R>0$, such that $T_{x,y}=0$ whenever $\mathrm{d}(x,y)>R$.
@@
    The expansion function of $f$, given by
    $\omega_{f}\colon[0,+\infty)\to[0,+\infty),\quad\omega_{f}(r)\mathrel{\vcentcolon=}\left\{\mathrm{d}_{Y}(f(x_{1}),f(x_{2}))\;\middle|\;\mathrm{d}_{X}(x_{1},x_{2})\leq r\right\},$
@@
$f_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X}))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(Y,\mathcal{H}_{Y}))$
@@
The coarse category consists of proper metric spaces as objects, and equivalence classes of coarse maps as arrows, where two coarse maps are equivalent iff they are close. A coarse equivalence is an isomorphism in the coarse category. If $X\subseteq Y$ is a metric subspace that is coarsely equivalent to $Y$, then we also say that $X$ is coarsely dense in $Y$. In particular, any Delone set $\Lambda$ of $\mathbb{R}^{d}$ is coarsely dense in $\mathbb{R}^{d}$.
@@
$\iota_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X}))\to\operatorname{KK}(\mathrm{C}\ell_{k,0},\mathrm{C}^{*}_{\mathrm{Roe}}(Y,\mathcal{H}_{Y})).$
@@
By 3.17, the induced maps of $\operatorname{Ad}_{V}$ coincides with that of $\operatorname{Ad}_{V^{\prime}}$, hence an isomorphism as well. ∎
@@
By Proposition 3.19, the Roe C*-algebra of a Delone set $\Lambda\subseteq\mathbb{R}^{d}$ is unique up to isomorphism, independent of its defining ample module. In particular, it is isomorphic to the Roe C*-algebra of $\mathbb{R}^{d}$. However, we must be careful as the inclusion $\iota\colon\Lambda\hookrightarrow\mathbb{R}^{d}$ does not induce an isomorphism (via a covering isometry) between the corresponding Roe C*-algebras $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda,\mathcal{H}_{\Lambda})$ and $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{R}^{d},\mathcal{H}_{\mathbb{R}^{d}})$ defined using their standard ample modules. An isomorphism between the Roe C*-algebras of coarsely equivalent spaces $X$ and $Y$ was constructed in [Ewert-Meyer:Coarse_geometry]*Theorem 3. However, these spaces carry different ample modules, involving an inverse of the coarsely dense embedding $\iota\colon X\to X\coprod Y$ in the coarse category.
As opposed to the groupoid C*-algebra $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ has simple K-theory groups. It follows from coarse Mayer–Vietoris sequence (cf. [Higson-Roe-Yu:Coarse_Mayer-Vietoris]) that the K-theory of $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{R}^{d})$ coincides with the K-theory of a point by a degree shift of $-d$. This holds for any Delone set $\Lambda\subseteq\mathbb{R}^{d}$ as well by Proposition 3.19. Therefore, K-theory of the real or complex Roe C*-algebras of a Delone set $\Lambda\subseteq\mathbb{R}^{d}$ is given by
(3.21) $\operatorname{K}_{i}(\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)_{\mathbb{C}})$$\simeq\begin{cases}\mathbb{Z}{\hphantom{/2}}&\text{if }i-d\equiv 0{\hphantom{,5,6,7}}\mkern-10.0mu\mod 2;\\ 0&\text{if }i-d\equiv 1{\hphantom{,5,6,7}}\mkern-10.0mu\mod 2.\end{cases}$
$\operatorname{K}_{i}(\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)_{\mathbb{R}})$$\simeq\begin{cases}\mathbb{Z}&\text{if }i-d\equiv 0,4{\hphantom{,6,7}}\mkern-10.0mu\mod 8;\\ \mathbb{Z}/2&\text{if }i-d\equiv 1,2{\hphantom{,6,7}}\mkern-10.0mu\mod 8;\\ 0&\text{if }i-d\equiv 3,5,6,7\mkern-10.0mu\mod 8.\end{cases}$
While using the Roe C*-algebra as the observable C*-algebra, there is no need to pass to its matrix algebras, as opposed to the case of groupoids. It was proven in [Ewert-Meyer:Coarse_geometry]*Corollary 1 that an operator $T$ on $\mathcal{H}_{\Lambda}^{\oplus N}$ is locally compact and controlled iff its matrix elements in $\mathbb{B}(\mathcal{H}_{\Lambda})$ are locally compact and controlled, when $T$ is viewed as an $N\times N$-matrix with entries in $\mathbb{B}(\mathcal{H}_{\Lambda})$. Therefore
@@
for all $N$, where $\mathcal{H}_{\Lambda}^{\oplus N}$ carries the direct sum of the standard ample representation $\mathrm{C}_{0}(\Lambda)\to\mathbb{B}(\mathcal{H}_{\Lambda})$. We may further identify $\mathcal{H}_{\Lambda}^{\oplus}\simeq\ell^{2}(\Lambda,\mathcal{K})\otimes\mathbb{R}^{N}\simeq\ell^{2}(\Lambda,\mathcal{K}\otimes\mathbb{R}^{N})\simeq\ell^{2}(\Lambda,\mathcal{K})$ using any unitary isomorphism $\mathcal{K}\simeq\mathcal{K}\otimes\mathbb{R}^{N}$. Such an isomorphism preserves controlled or locally compact operators using the characterisation in Example 3.15.

### v2

@@ 3.1.1. The regular representation
The reduced real C*-algebra of an étale groupoid $\mathcal{G}$ is the completion of the real groupoid convolution algebra $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$ under the regular representation (cf. [Khoshkam-Skandalis:Regular_representation]), which we recall here. The real groupoid convolution algebra $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$ consists of real-valued, compactly supported continuous functions $f\colon\mathcal{G}\to\mathbb{R}$, equipped with the convolution product and involution
@@
The space $\mathrm{C}_{\mathrm{c}}(\mathcal{G})$ is a pre-Hilbert module over $\mathrm{C}_{0}(\mathcal{G}^{0})$, with right $\mathrm{C}_{0}(\mathcal{G}^{0})$-module structure and inner product (cf. [Bourne-Mesland:Topological_phases]*Section 1.3)
$(f\cdot\varphi)(\gamma)$ $\mathrel{\vcentcolon=}f(\gamma)\varphi(s(\gamma)),$
$\left\langle f,g\right\rangle(x)$ $\mathrel{\vcentcolon=}(f^{*}*g)\big|_{\mathcal{G}^{0}}(x)=\sum_{\gamma\in r^{-1}(x)}f(\gamma)g(\gamma),$
@@
$\left\{(\omega-x,-x)\;\middle|\;0\in\omega-x\right\}=\left\{(\omega-x,-x)\;\middle|\;x\in\omega\right\},$
which is in bijection with the Delone set $\omega$ via $(\omega-x,-x)\mapsto x$. Thus the localised Hilbert space $\mathcal{H}_{\omega}$ is unitarily isomorphic to $\ell^{2}(\omega)$ via (cf. [Bourne-Mesland:Topological_phases]*Section 4.1):
@@
Lemma 3.8 ([Bourne-Mesland:Topological_phases]*Section 4.1).
@@
When we consider systems with a finite number of internal degrees of freedom inside every lattice site, e.g. spins of electrons, then we must replace the observable C*-algebra by its matrix algebra $\mathbb{M}_{N}\mathrm{C}^{*}(\mathcal{G}_{\Lambda})=\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})$. In the same way as Section˜2.3, we extend the regular representation of $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ to a representation
$\pi_{\omega}^{N}\mathrel{\vcentcolon=}\pi_{\omega}\otimes e_{N},\quad\mathrm{C}^{*}(\mathcal{G}_{\Lambda})\otimes\mathbb{M}_{N}(\mathbb{R})\to\mathbb{B}(\ell^{2}(\omega)\otimes\mathcal{K}),$
where $\mathcal{K}$ is a separable Hilbert space, and $e_{N}\colon\mathbb{M}_{N}(\mathbb{R})\hookrightarrow\mathbb{K}(\mathcal{K})$ is any rank-$N$ corner embedding induced by an isometry $\mathbb{R}^{N}\hookrightarrow\mathcal{K}$. We may describe the representation $\pi_{\omega}^{N}$ by an infinite matrix indexed by $x,y\in\omega$, with matrix elements
@@ 3.1.2. Position spectral triple over the groupoid C*-algebra
In order to construct a spectral triple over $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$ or $\mathbb{M}_{N}\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, one way is to first construct an unbounded Kasparov $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$-$\mathrm{C}(\Omega_{0})$–module, then take its (unbounded) Kasparov product with a *-homomorphism $\mathrm{C}(\Omega_{0})\to\mathbb{R}$. Such an unbounded Kasparov module was given in [Bourne-Mesland:Topological_phases]*Section 2.3.1 using the “position” operators on the Hilbert $\mathrm{C}(\Omega_{0})$-module $L^{2}(\mathcal{G}_{\Lambda})$:
@@
The map $c_{j}\colon\mathcal{G}_{\Lambda}\to\mathbb{R}$ is an exact groupoid cocycle in the sense of [Mesland:Groupoid_cocycles]*Definition 4.1.2. It follows from the general construction in [Mesland:Groupoid_cocycles]*Theorem 3.2.2 that $\prescript{}{d}{\lambda}_{\Omega_{0}}$ is an unbounded Kasparov module. Following [Bourne-Mesland:Topological_phases], we call $\prescript{}{d}{\lambda}_{\Omega_{0}}$ the bulk cycle.
Now we construct the position spectral triple by localising $\prescript{}{d}{\lambda}_{\Omega_{0}}$ at any $\omega\in\Omega_{0}$. As explained in the paragraph after Definition˜3.1, the physically relevant Hilbert space is $\ell^{2}(\omega)\otimes\mathcal{K}$ instead of $\ell^{2}(\omega)$ or $\ell^{2}(\omega)\otimes\mathbb{R}^{N}$. This is fixed by taking any isometry $\mathbb{R}^{N}\hookrightarrow\mathcal{K}$ and its induced corner embedding $e_{N}\colon\mathbb{M}_{N}(\mathbb{R})\hookrightarrow\mathbb{K}(\mathcal{K})$.
@@
(3.13) $\xi_{\omega,N}^{\mathrm{Gpd}}\mathrel{\vcentcolon=}\left(\mathbb{M}_{N}(\mathrm{C}_{\mathrm{c}}(\mathcal{G}_{\Lambda}))\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\omega)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right)$
is a position spectral triple. It represents a class in
@@
which is the Kasparov product of the following (unbounded) Kasparov modules and *-homomorphisms:
@@
    the corner embedding $e_{N}\colon\mathbb{M}_{N}(\mathbb{R})\hookrightarrow\mathbb{K}(\mathcal{K})$, which gives an invertible element in $\operatorname{KK}(\mathbb{M}_{N}(\mathbb{R}),\mathbb{R})$.
@@
is isomorphic to $\ell^{2}(\omega)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$ via the localised representation at $\omega$. Thus it suffices to check that $c_{j}$ is mapped to ${\scriptstyle\mathsf{X}}_{j}$ under the composition and that (3.13) is a spectral triple. By (3.7), the operator $X_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\operatorname{id}_{\bigwedge\nolimits^{*}\mathbb{R}^{d}}$ acts on $\ell^{2}(\omega)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d}$ sends $\left\lvert x\right\rangle\otimes v\otimes w$ to $x_{j}\left\lvert x\right\rangle\otimes v\otimes w$, where $x\in\omega$ is considered an element in $\mathbb{R}^{d}$. It is a spectral triple by [Bourne-Mesland:Topological_phases]*Proposition 4.1. Thus (3.13) is a position spectral triple. ∎
@@ 3.2. Roe C*-algebra of a Delone set
An ample $X$-module or ample module over $X$ is given by a pair $(X,\mathcal{H}_{X})$, where $X$ is a proper metric space, and $\mathcal{H}_{X}$ is a Hilbert space that carries a non-degenerate representation $\varrho\colon\mathrm{C}_{0}(X)\to\mathbb{B}(\mathcal{H}_{X})$, which is ample in the sense that
$\varrho(f)\in\mathbb{K}(\mathcal{H}_{X})$ if and only if $f=0$.
@@
  • An operator $T\in\mathbb{B}(\mathcal{H}_{X})$ is called locally compact, if for any $f\in\mathrm{C}_{\mathrm{c}}(X)$, $T\varrho(f)$ and $\varrho(f)T$ are compact operators.
  • An operator $T\in\mathbb{B}(\mathcal{H}_{X})$ is called controlled, if $\operatorname{supp}(T)$ is controlled.
@@
Let $\Lambda$ be a discrete metric space and $\mathcal{H}_{\Lambda}=\ell^{2}(\Lambda,\mathcal{K})$ be the standard ample $\Lambda$-module. Describe an operator $T\in\mathbb{B}(\mathcal{H}_{\Lambda})$ by an infinite matrix $(T_{x,y})_{x,y\in\Lambda}$ where
@@
The operator $T$ is locally compact, if and only if $T_{x,y}\in\mathbb{K}(\mathcal{K})$ for all $x,y\in\Lambda$; controlled, if and only if there exists $R>0$, such that $T_{x,y}=0$ whenever $\mathrm{d}(x,y)>R$.
@@
    The expansion function $\omega_{f}\colon[0,\infty)\to[0,\infty)$, given by the formula
    $\omega_{f}(r)\mathrel{\vcentcolon=}\sup\left\{\mathrm{d}_{Y}(f(x_{1}),f(x_{2}))\;\middle|\;\mathrm{d}_{X}(x_{1},x_{2})\leq r\right\}$
@@
$f_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X}))\to\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(Y,\mathcal{H}_{Y}))$
@@
The coarse category consists of proper metric spaces as objects, and equivalence classes of coarse maps as arrows, where two coarse maps are equivalent if and only if they are close. A coarse equivalence is an isomorphism in the coarse category. If $X\subseteq Y$ is a metric subspace that is coarsely equivalent to $Y$, then we also say that $X$ is coarsely dense in $Y$. In particular, any Delone set $\Lambda$ of $\mathbb{R}^{d}$ is coarsely dense in $\mathbb{R}^{d}$.
@@
$\iota_{*}\colon\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X}))\to\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(Y,\mathcal{H}_{Y})).$
@@
By ˜3.20, the induced maps of $\operatorname{Ad}_{V}$ coincides with that of $\operatorname{Ad}_{V^{\prime}}$, hence an isomorphism as well. ∎
@@
By Propositions˜3.21 and 3.22, given two Delone set $\Lambda_{1},\Lambda_{2}\subseteq\mathbb{R}^{d}$ and ample modules $\mathcal{H}_{\Lambda_{1}},\mathcal{H}_{\Lambda_{2}}$, then there exists an isomorphism $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda_{1},\mathcal{H}_{\Lambda_{1}})\simeq\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda_{2},\mathcal{H}_{\Lambda_{2}})$. More generally, all of them are isomorphic to the Roe C*-algebra of $\mathbb{R}^{d}$ defined by its standard ample module, cf. [Ewert-Meyer:Coarse_geometry]*Theorem 2. However, we note that the inclusion $\iota\colon\Lambda\hookrightarrow\mathbb{R}^{d}$ does not induce an isomorphism (via a covering isometry) between the corresponding Roe C*-algebras $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda,\mathcal{H}_{\Lambda})$ and $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{R}^{d},\mathcal{H}_{\mathbb{R}^{d}})$ defined using their standard ample modules. An isomorphism between the Roe C*-algebras of coarsely equivalent spaces $X$ and $Y$ was constructed in [Ewert-Meyer:Coarse_geometry]*Theorem 3. However, these spaces carry different ample modules, involving an inverse of the coarsely dense embedding $\iota\colon X\to X\coprod Y$ in the coarse category.
As opposed to the groupoid C*-algebra $\mathrm{C}^{*}(\mathcal{G}_{\Lambda})$, the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$ has simple K-theory groups. It follows from coarse Mayer–Vietoris sequence (cf. [Higson-Roe-Yu:Coarse_Mayer-Vietoris]) that the K-theory of $\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{R}^{d})$ coincides with the K-theory of a point by a degree shift of d-d. This holds for any Delone set $\Lambda\subseteq\mathbb{R}^{d}$ as well by Proposition˜3.22. Therefore, K-theory of the real or complex Roe C*-algebras of a Delone set $\Lambda\subseteq\mathbb{R}^{d}$ is given by
[…]
We note that the isomorphisms of abelian groups in (3.24) are actually given by $\operatorname{K}_{*}(\mathbb{C})$-module or $\operatorname{KO}_{*}(\mathbb{R})$-module isomorphisms
$\operatorname{K}_{*}(\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)_{\mathbb{C}})\simeq\operatorname{K}_{*-d}(\mathbb{C}),\quad\operatorname{KO}_{*}(\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)_{\mathbb{R}})\simeq\operatorname{KO}_{*-d}(\mathbb{R}),$
cf. [Ewert-Meyer:Coarse_geometry]*Section 3 and Corollary 5.
While using the Roe C*-algebra as the observable C*-algebra, there is no need to pass to its matrix algebras, as opposed to the case of groupoids. It was proven in [Ewert-Meyer:Coarse_geometry]*Corollary 1 that an operator $T$ on $\mathcal{H}_{\Lambda}^{\oplus N}$ is locally compact and controlled if and only if its matrix elements in $\mathbb{B}(\mathcal{H}_{\Lambda})$ are locally compact and controlled, when $T$ is viewed as an $N\times N$-matrix with entries in $\mathbb{B}(\mathcal{H}_{\Lambda})$. Therefore
@@
for all $N$, where $\mathcal{H}_{\Lambda}^{\oplus N}$ carries the direct sum of the standard ample representation $\mathrm{C}_{0}(\Lambda)\to\mathbb{B}(\mathcal{H}_{\Lambda})$. We may further identify $\mathcal{H}_{\Lambda}^{\oplus}\simeq\ell^{2}(\Lambda,\mathcal{K})\otimes\mathbb{R}^{N}\simeq\ell^{2}(\Lambda,\mathcal{K}\otimes\mathbb{R}^{N})\simeq\ell^{2}(\Lambda,\mathcal{K})$ using any unitary isomorphism $\mathcal{K}\simeq\mathcal{K}\otimes\mathbb{R}^{N}$. Such an isomorphism preserves controlled or locally compact operators using the characterisation in Example˜3.18.